OpenStudy (lilsis76):

How can I re-write the sum's suing Sigma Notation for: 2 - 2^2 + 2^3 -2^4 +2^5 I know the reoccuring number is going to be 2 which goes to the right of the sigma. on the bottom of sigma it would be 1, and top of sigma would be 5 right?

OpenStudy (lilsis76):



OpenStudy (anonymous):

\[\sum_{n=0}^{4}2(-2)^n\] is one way of writing it.

OpenStudy (lilsis76):

oh. so that (-2), is that how it would go from negative to positive?

OpenStudy (anonymous):

expand what i wrote and check for yourself

OpenStudy (lilsis76):

okay: -2^0 = 0 *2 = 0 -2^1=-2*2=-4 -2^2=-4*2=-8 -2^3=-8*2=-16

OpenStudy (lilsis76):

@cnknd is this how it would go?

OpenStudy (anonymous):

ur doing the exponents wrong... this is basics...

OpenStudy (lilsis76):

okay the first one would be>....< @cnknd